# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ Appendix E Proofs of Supporting Lemmas
E.9 Proof of Lemma D.7
Proof of Lemma D.7.
By Lemma D.6,
@@
Here, for the last inequality, we apply $\sum_{i=1}^{n}w_{i}=1$ and $\frac{{\rm d}\mu_{Z}}{{\rm d}\mu_{Z|W=0}}(z)\geq\mathbb{P}(W=0)~{}\forall z$, where $\mathbb{P}(W=0)=\int\mathbb{P}(W=0|Z=z){\rm d}\mu_{Z}(z)\geq\delta$ by Assp. 2.3.
@@
For (Term B), we apply Assp. 4.6 and get
$\mathbb{E}[(\text{Term B})]\leq\frac{\sqrt{d_{Z}}C_{w}}{2\delta}n^{-r}.$
@@
E.10 Proof of Lemma D.8
Proof of Lemma D.8.
We write $Z_{i}(0)$ as $Z_{i}$ in this proof. By the triangle inequality (Lemma A.1),
@@
For (Term A), apply Lemma D.7, we get
[…]
For (Term B), we follow the similar argument of the proof of [fournier2015rate, Theorem 1], and see that it depends on the bound of $\mathbb{E}[|\tilde{\mu}_{Z(0),n}(A)-\mu_{Z(0)}(A)|]$ for any subset $A$ of $\mathbb{R}^{d_{Z}}$.
@@
[…]
@@
Further, $\mathbb{P}(W=0)=\int\mathbb{P}(W=0|Z=z){\rm d}\mu_{Z}(z)\geq\delta$ by Assp. 2.3.
@@
This inequality is similar to the first inequality in [fournier2015rate, Section 3], except that instead of $\delta^{-2}$, they have $1$ as the constant.
Then, following the same argument as in the proof of [fournier2015rate, Theorem 1], we get
[…]
@@
E.11 Proof of Lemma D.9
Proof of Lemma D.9.
@@
then by the triangle inequality (Lemma A.1), we have
@@
where the last inequality is due to Lemma D.2.
For (Term A), by Lemma D.6, we have
[…]
@@
For $(\text{Term C})_{i}$, by the same reasoning above, we have

### v2

@@ H.9 Proof of Lemma 11
H.10 Proof of Lemma 12
Proof of Lemma 12.
By Lemma 11,
@@ H.10 Proof of Lemma 12
Here, for the last inequality, we apply $\sum_{i=1}^{n}w_{i}=1$ and $\frac{{\rm d}\mu_{Z}}{{\rm d}\mu_{Z|W=0}}(z)\geq\mathbb{P}(W=0)~\forall z$, where $\mathbb{P}(W=0)=\int\mathbb{P}(W=0|Z=z){\rm d}\mu_{Z}(z)\geq\delta$ by Assp. 3.
@@
For (Term B), we apply Assp. 13 and get
$\mathbb{E}[(\text{Term B})]\leq\frac{\sqrt{d_{Z}}C_{\textup{w}}}{2\delta}n^{-r}.$
@@
H.11 Proof of Lemma 13
Proof of Lemma 13.
We write $Z_{i}(0)$ as $Z_{i}$ in this proof. By the triangle inequality (Lemma 1),
@@ H.11 Proof of Lemma 13
For (Term A), apply Lemma 12, we get
$\text{(Term A)}\leq\frac{\sqrt{d_{Z}}}{2\delta^{3}}n^{-\frac{1}{2}}+\frac{\sqrt{d_{Z}}C_{\textup{w}}}{2\delta}n^{-r}.$
For (Term B), we follow the similar argument of the proof of [23, Theorem 1], and see that it depends on the bound of $\mathbb{E}[|\tilde{\mu}_{Z(0),n}(A)-\mu_{Z(0)}(A)|]$ for any subset $A$ of $\mathbb{R}^{d_{Z}}$.
@@
$\leq$ $\frac{1}{\sqrt{n}\mathbb{P}(W=0)}\left(\mathbb{E}_{Z\sim\mu}\left[\left(\frac{{\rm d}\mu_{Z}}{{\rm d}\mu_{Z|W=0}}(Z)\delta_{Z}(A)-\mu_{Z(0)}(A)\right)^{2}\right]\right)^{\frac{1}{2}}$ $(\text{Jensen's inequality})$
@@
Further, $\mathbb{P}(W=0)=\int\mathbb{P}(W=0|Z=z){\rm d}\mu_{Z}(z)\geq\delta$ by Assp. 3.
@@
This inequality is similar to the first inequality in [23, Section 3], except that instead of $\delta^{-2}$, they have $1$ as the constant.
Then, following the same argument as in the proof of [23, Theorem 1], we get
$\text{(Term B)}\leq\mathbb{E}[\mathds{W}_{1}(\tilde{\nu},\nu)]\leq C\delta^{-2}R_{d_{Z}}(n),$
@@
H.12 Proof of Lemma 14
Proof of Lemma 14.
@@ H.12 Proof of Lemma 14
then by the triangle inequality (Lemma 1), we have
@@
where the last inequality is due to Lemma 7.
For (Term A), by Lemma 11, we have
$(\text{Term A})\leq$ $\frac{\sqrt{d_{Z}}}{2}\sum_{G\in\Phi_{r}^{n}}\sum_{i\in G}\left|\frac{W_{G}}{|\{i:Z_{i}(0)\in G\}|}-\hat{w}_{i}\right|$
@@
[…]